How can I write these coordinates, (2,-3) and (-3,7) in slope intercept form, standard form, and point-intercept form?
Answer by rfer(16322)

slope m=(7--3)/(-3-2)
m=10/-5
m=-2
----------------
y=mx+b
7=-2*-3+b
7=6+b
7-6=b
1=b
==============
y=-2x+1
you do the others.